# SPLITZ: Certifiable Robustness via Split Lipschitz Randomized Smoothing thanks: This work was supported by NSF grants CCF 2100013, CNS 2209951, CCF 1651492, CNS 2317192 and CNS 1822071. This work has been submitted to IEEE Transactions on Information Forensics and Security

### v2

@@ V Discussion and conclusion
In this paper, we presented SPLITZ, a novel and practical certified defense mechanism, where we constrained the local Lipschitz bound of the left half of the classifier and smoothed the right half of the classifier with noise. This is because the local Lipschitz constant can capture information specific to each individual input, and the relative stability of the model around that input. To the best of our knowledge, this is the first systematic framework to combine the ideas of local Lipschitz constants with randomized smoothing. Furthermore, we provide a closed-form expression for the certified radius based on the local Lipschitz constant of the left half of the classifier and the randomized smoothing based radius of the right half of the classifier. We show that maintaining a relatively small local Lipschitz constant of the left half of the classifier helps to improve the certified robustness (radius). We showed results on several benchmark datasets and obtained significant improvements over state-of-the-art methods for the MNIST, CIFAR-10 and ImageNet datasets. For instance, on the CIFAR-10 dataset, SPLITZ can achieve $43.2\%$ certified test accuracy compared to state-of-art certified test accuracy $39.8\%$ with $\ell_{2}$ norm perturbation budget of $\epsilon=1$. We observed similar trends for the MNIST and ImageNet dataset. We believe that combining the core idea of SPLITZ with other recent techniques, such as denoising diffusion models and adversarial training can be a fruitful next step to further improve certified robustness.

### v3

@@ V Discussion and conclusion
[…]
@@ Appendix C Additional Experimental Details
MNIST Certified Test Accuracy at $\epsilon~(\%)$
Method Extra data 1.50 1.75 2.00 2.25 2.50
RS [7] 67.3 46.2 32.5 19.7 10.9
MACER [19] 73.0 50.0 36.0 28.0 -
Consistency [25] 82.2 70.5 45.5 37.2 28.0
SmoothMix [15] 81.8 70.7 44.9 37.1 29.3
DRT [26] 83.3 69.6 48.3 40.3 34.8
RS+OrthoNN [24] 70.1 49.7 33.2 21.0 10.5
SPLITZ 80.2 71.3 62.3 51.7 40.5
TABLE X: Comparison of certified test accuracy (%) on MNIST under $\ell_{2}$ norm perturbation. Each entry lists the certified accuracy using numbers taken from respective papers (RS results follow from previous benchmark papers [25, 15]).
Results on MNIST As showed in Table X, we can observe that SPLITZ outperforms other state-of-art approaches in almost every value of $\epsilon$. Impressively, we find that the SPLITZ classifier has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=2.50$, SPLITZ classifier achieves 40.5% compared to state-of-art top-1 test accuracy 34.8% certified test accuracy on the MNIST dataset. Moreover, when we increase $\epsilon$ from 1.50 to 2.50, RS drops from 67.3% to 10.9% decreasing 56.4% test accuracy. SPLITZ, however, maintains higher certified test accuracy from 80.2% to 40.5% maintaining relatively higher test accuracy.